OpenStudy (anonymous):

3^-2x+1 times 3^-2x-3=3^-x

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[3^{-2x+1}\times 3^{-2x-3}=3^{-x}\]?

OpenStudy (anonymous):

yes it is. im very confused.

OpenStudy (misty1212):

ok we can do this, it is not that hard

OpenStudy (misty1212):

when you multiply you add the exponents

OpenStudy (misty1212):

therefore the left hand side can be written as \[\huge 3^{-2x+1-2x-3}\] combine like terms, get \[\huge 3^{-4x-2}\]

OpenStudy (misty1212):

since now the base on both sides is 3, the exponents must be equal as well your final job is to solve \[-4x-2=-x\] for \(x\) which is routine i imagine

OpenStudy (misty1212):

you good from there?

OpenStudy (anonymous):

I think so, thank you so much!(:

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

ill just tell you the answer hold on a sec

OpenStudy (anonymous):

not sure if your solving for x but if you are than the answer is x=2

OpenStudy (anonymous):

are you here to waste my time or to tell me if its solving for x or not